# thanks: Deceasedthanks: Deceasedthanks: Deceased

### v1

@@ IV In-flight experimental campaign
At the beginning of the commissioning period, all subsystems went through an initial checkout procedure. In this initial phase, coil #$\#2$ —the one closest to TM2— showed a malfunctioning during the verification. Due to this fact, the injections performed during the operations period were on coil #$\#1$, and the only set of injections performed in coil #$\#2$ were done with low currents to prevent any possible current leak to other systems, which resulted in a reduction of the precision achievable with coil #$\#2$ experiments. This implies that most of the results that will be shown here will be for TM1 if not specified otherwise.
@@
We carried a total of three sets of magnetics injections. The first set was injected on the days $\rm 28^{th}$ and $\rm 29^{th}$ of April, 2016. The injection of day $\rm 28^{th}$ consisted on applying a sinusoidal signal through the coil #$\#1$ at different DCs and with different AC amplitudes. The injections on day $\rm 29^{th}$ were exactly the same but through coil #$\#2$. The second set of injections were carried on the $\rm 18^{th}$ of June, 2016. It consisted on a series of sinusoidal injections at a wider range of both DCs and ACs than the previous ones and exclusively in coil #$\#1$. The third, and last, set of injections were performed on the days $\rm 14^{th}$, $\rm 15^{th}$ and $\rm 16^{th}$ of March, 2017. They consisted on a very long lasting signal at a high DC and with a small sinusoidal on top of it. These were applied exclusively through coil #$\#1$ too. The complete list of experiments is shown in Appendix C.
@@
A typical run of magnetic experiments is shown in Figure 2. The three panels display the main variables of interest in our analysis, these are the magnetic field in the x direction, as measured by the closest magnetometer to each coil, the acceleration produced between the TMs due to the presence of these injections and the torque being induced between both TMs along the y axis.
We evaluate the magnetically induced force in the test mass through the $\Delta g$ variable —the end scientific output of the mission— nominally defined as the differential stray acceleration between the two TMs in their nominal position [20]. Since the main objective of the $\Delta g$ is the evaluation of the free fall of the test masses, those forces arising in the spacecraft dynamics control loop or others forces originating in the LISA Pathfinder non-inertial reference frame are rightfully subtracted in the definition of this parameter [21]. Indeed, the terms $\hat{F}_{DC,x}$, $\hat{F}_{1\omega,x}$ and $\hat{F}_{2\omega,x}$ previously defined in Section III.2 are estimated by demodulating the $\Delta g$ at the corresponding frequencies and rescaling the amplitudes obtained by means of the mass of the TMs, $\rm{m_{TM}}$ = 1.928 kg. Analogously, the same procedure can be extrapolated to the torque by using the moment of inertia of a cube with the side length of the TMs, 46 mm.
@@
Since the magnetic field can not be directly measured in the test mass position we must refer to the magnetometers read-out for calibration. Hence, we estimated the amplitude in the PX magnetometer of each of the 20 injections of June 18th, 2016 for coil #$\#1$. To do so we demodulate the measured read-out at the injection frequency. By comparing these amplitudes to the ones predicted by Ampère’s law we found a systematic discrepancy of $11.85\pm 0.45\%$ being the predicted magnetic field larger than the measurements from the magnetometers. Magnetometers accuracy of 0.5% allows the most precise calibration to prevent any mechanical tolerances during manufacturing and assembly of the coils and displacements and tilts that may have arisen during launch. In the following, we will apply this correction to all the calculated magnetic field values that intervene in all the equations derived in the previous section and redefine them for simplicity, i.e. $B_{x}^{AC}\equiv 0.8815\left\langle B_{x}^{AC}\right\rangle$.
@@ IV.1 Remanent magnetic moment
$\hat{F}_{1\omega,x}=\left[M_{x}+\frac{2\chi V}{\mu_{0}}B_{x}^{DC}\right]\nabla_{x}B_{x}^{AC}.$ (14)
@@
In order to evaluate this term, during the June $\rm 18^{th}$ run several injections with different AC field gradients were applied to the TMs. By doing so we can evaluate the term in brackets above at different values of the gradient $\nabla_{x}B_{x}^{AC}$. This is shown in Figure 3 where we display how the $1\omega$ component of the force changes depending on the intensity of the injected AC magnetic field gradient, for fixed DCs. Each result in the plot represents an injection at 5 mHz with AC amplitudes applied to the coil of $I^{AC}=0.5,0.8,1.0,1.5\,{\rm mA}$.
As explained in Section III.2, by running the experiment at different DC levels we can further disentangle the dependencies for the parameters inside the brackets and obtain the estimate for the remanent magnetic moment. Table 1 gathers the linear fits to the results that we also show in Figure 4. The offset parameter corresponds to the remanent magnetic moment of test mass #$\#1$ in the x direction, $M_{x}=0.140\pm 0.138$ nAm2. The slope parameter is directly related to the magnetic susceptibility at $1\omega$. However, the values that were used for DCs of ±$\pm 0.1,0.2$ mA came from the injections of the $\rm 28^{th}$ of April which were performed at different frequencies, 3 mHz, than the rest at 5 mHz. Thus, to obtain the value of the susceptibility at 5 mHz, we will use the fit from Fig. 4, but with only the DC values of ±$\pm 0.75,1.5$ mA giving a result of $\chi_{\rm{5mHz}}=(-3.3723\pm 0.0069)\times 10^{-5}$.
@@
Figure 4: Effective remanent magnetic moment plotted as a function of the injected DC magnetic field. The result of the fit, for an equation of the type y = mx + n, is m = $(-3.380\pm 0.027)\times 10^{-5}$ and n = $(0.142\pm 0.140)\times 10^{-9}$ Am2.
Analogously, to obtain $M_{y}$ and $M_{z}$, we demodulate the amplitudes of the torque measurements around the required angles and apply Eqs. (12) for the respective injected magnetic fields. This way, we obtained $M_{y}=0.178\pm 0.025$ nAm2 and $M_{z}=0.095\pm 0.010$ nAm2 which we will need for the background estimations. These results lead to a total remanent magnetic moment of: $|\textbf{M}|=(0.245\pm 0.081)$ nAm2.
@@ IV.2 Background magnetic field
The induction of forces in the test mass by means of the controlled injection of magnetic fields allow the determination not only of the test mass magnetic parameters but also of environment parameters that contribute to the magnetic force, which is the case of the background magnetic field in the test mass position. We emphasize here that this is the only precise estimate of this parameter since the magnetometers are located too far away from the test masses to guarantee a precise estimate of this variable.
In order to do so we evaluate Eqs. (13) using the injections of June 18th, 2016. We express the information provided by these runs by displaying the DC component of the measured force as a function of the DC component of the applied magnetic fields. As predicted, we obtain the parabolas in Fig. 5 for different values of the AC amplitude from where we derive the parabola coefficients of Table 2.
@@
As derived from Eqs. (13), the A coefficient provides a direct estimate of the magnetic susceptibility
@@
which, in comparison with other alternative estimates, is not dependent of the injection modulation frequency. The value obtained using equation (15) is $\chi_{DC}=(-3.35\pm 0.15)\times 10^{-5}$.
With the two remaining terms, B and C, we can build a system of equations, being the two unknowns the parameters that define the background magnetic field in the test mass position, i.e. $B_{back.,x}$ and $\nabla_{x}B_{back.,x}$
@@
$B=\frac{M_{x}}{\kappa}+\frac{\chi V}{\mu_{0}}\left[\nabla_{x}B_{back.,x}+\frac{B_{back.,x}}{\kappa}\right].$ (17)
@@ IV.3 Magnetic Susceptibility
We have already estimated the test mass magnetic susceptibility as a by product of the estimate of the remanent magnetic moment and the background magnetic field and field gradient.
@@
If we now plot the measured values of the magnetic susceptibility for TMs 1 and 2 along with the curve in Eq. (18), we obtain the plot shown in Figure 6.
The results shown were obtained by heterodyning the $\Delta g$ signal with a sinusoidal signal in phase. However, if we made the latter be out of phase by $\pi/2$ one would expect the amplitudes measured by this method to be zero only if there was no imaginary component. Thus, the imaginary susceptibility can be obtained demodulating at $2\omega$, in quadrature. At the frequency regime that we are working on, around the mHz, this value is expected to be orders of magnitude smaller than the real susceptibility. The results that we obtain for the imaginary susceptibility at 10 mHz were consistent with zero, $\chi_{i}=(0.0\pm 1.8)\times 10^{-6}$. This is because the method employed to measure the amplitudes of the $\Delta g$ was limited around the tenths of femtonewtons. So, we can only confirm that the values of the imaginary susceptibility of the TMs are below $|\chi_{i}|<1.8\times 10^{-6}$ at 10 mHz. At the rest of frequencies the result gave a less precise upper bound for the imaginary susceptibility.
Finally, once all parameters are determined, we can evaluate the prediction of our force model in comparison with the measured acceleration during injections. To do so, we have selected a single injection from all the ones of June $18^{th}$, 2016 with amplitudes of current DC = 0.75 mA and AC = 1.5 mA. We have used these values for the corresponding magnetic fields DC and AC calculations respectively together with the magnetic parameters found in the previous sections into Eqs. (7). The results can be seen in Fig. 7 where, together with the total model force, we can also see plotted the different force contributions $F_{DC}$, $F_{1\omega}$ and $F_{2\omega}$. The predicted force and the data match with a residual difference between them of $(0.0\pm 1.9)\times 10^{-13}$ N.
@@ V Conclusions
The work in this paper presents the most detailed characterisation of the magnetic-induced coupling on free-falling test masses in the context of gravitational wave detection in space. The results and method showed here were originated in the framework of the LISA Pathfinder mission. As a technology demonstrator, LPF represented a unique opportunity for an in-depth description of these effects, which can be directly transferred to LISA and future space-borne gravitational wave detectors.
In our results, we obtain the magnetic parameters defining the response of the free-falling test masses under an external magnetic field. According to these, the test mass remanent magnetic moment do not show any privileged direction, which is in agreement with the hypothesis of an isotropic diamagnetic test mass. We have found the value of its modulus to be: $\rm|\textbf{M}|=(0.245\pm 0.081)\,\rm{nAm}^{2}$, a result with much better precision than any other on-ground tests and also below the mission requirements of $\rm|\textbf{M}|\leq 10\,nAm^{2}$. The background magnetic field, $B_{back.,x}=414\pm 74$ nT, and its gradient, $\nabla_{x}B_{back.,x}=-7400\pm 2100$ nT/m, calculated at the location of the TMs are in agreement with the values predicted by  [22]. Moreover, the gradient is also below the worst case scenario predicted by  [25] which estimated a value of $\nabla_{x}B_{back.,x}=-11300$ nT/m. The TMs magnetic susceptibility results follow the model predicted by  [23] and they are also within the same magnitude when comparing the on-ground estimations for the DC magnetic susceptibility with values around $-2.5\times 10^{-5}$  [26] with our in-flight DC prediction for the magnetic susceptibility: $\chi_{DC}=(-3.35\pm 0.15)\times 10^{-5}$. Finally, the value of the imaginary component of the magnetic susceptibility of the TMs has been set to an upper bound at 10 mHz such that: $|\chi_{i}|<1.8\times 10^{-6}$.
These results allow the sound design of future gravitational detectors in space based on results obtained in-flight. They are also key in either reducing or suppressing the need for such a characterisation for future missions, which would otherwise need dedicated time during its commissioning period.
@@ Appendix A Magnetic field calculation
When the magnetic field is created by an induction coil, such as the case in Fig 1, the laws of Classical Magnetic Theory can be applied to obtain formulas which produce the values of the field and its gradient at any position in space. For slowly varying coil currents and short distances, radiative effects can be neglected.
@@ Appendix B Magnetic torque
$\begin{split}\text{N}_{\eta}=&\left\langle\sin(\omega t)\left(m_{z}B_{x}^{AC}-m_{x}B_{z}^{AC}\right)\right\rangle\\ &+\left\langle\left[r_{z}\Omega_{x}-r_{x}\Omega_{z}\right]\right\rangle,\end{split}$ (44)
@@
$\begin{split}\mathbf{\Omega}=&\left(\mathbf{m}\cdot\mathbf{\nabla}\right)\mathbf{B}^{AC}\sin(\omega t)\\ &+\dfrac{\chi}{\mu_{0}}\left(\dfrac{1-\cos(2\omega t)}{2}\right)\left(\mathbf{B}^{AC}\mathbf{\nabla}\right)\mathbf{B}^{AC},\end{split}$ (45)
@@
$\text{N}_{\eta}=\left\langle m_{z}B_{x}^{AC}\right\rangle\sin(\omega\,t).$ (46)
@@
$\text{N}_{\phi}=\left\langle-m_{y}B_{x}^{AC}\right\rangle\sin(\omega\,t).$ (47)
@@
$\begin{split}\mathbf{N}_{DC}=\left\langle\mathbf{m}\times\mathbf{B}_{0}\right\rangle,\end{split}$ (49a)
$\begin{split}\mathbf{N}_{1\omega}=\left\langle\mathbf{m}\times\mathbf{B}^{AC}\right\rangle\sin(\omega\,t).\end{split}$ (49b)

### v2

@@ IV In-flight experimental campaign
At the beginning of the commissioning period, all subsystems went through an initial checkout procedure. In this initial phase, coil #$\#2$ —the one closest to TM2— showed a malfunction. Due to this fact, the injections performed during the operations period were on coil #$\#1$, and the only set of injections performed in coil #$\#2$ were done with low currents to prevent any possible current leak to other systems. This resulted in a reduction of the precision achievable with coil #$\#2$ experiments. Thus, most of the results that will be shown here will be for TM1 if not specified otherwise.
@@
We carried out a total of three sets of magnetics injections. The first set was performed on the April $\rm 28^{th}$ and $\rm 29^{th}$, 2016. The injections of April $\rm 28^{th}$ consisted of applying a sinusoidal signal through the coil #$\#1$ with different DC offsets and different AC amplitudes. The injections of April $\rm 29^{th}$ were exactly the same but through coil #$\#2$. The second set of injections were done on June $\rm 18^{th}$, 2016. They consisted of a series of sinusoidal injections at a wider range of both DC offsets and AC amplitudes than the previous ones but performed exclusively in coil #$\#1$. The third, and last, set of injections were performed on March $\rm 14^{th}$, $\rm 15^{th}$ and $\rm 16^{th}$ , 2017. They consisted of very long-lasting signals at high DC offsets and with small AC amplitudes applied exclusively through coil #$\#1$. The complete list of experiments is shown in Appendix C.
@@
A typical run of magnetic experiments is shown in Fig. 2, where we show the results from the third set of injections during June $\rm 18^{th}$. The applied currents through the coil at 5 mHz can be seen in the Appendix within Table 6. The three panels display the main variables of interest in our analysis, these are the magnetic field in the x direction, as measured by the closest magnetometer to each coil, the acceleration produced between the TMs due to the presence of these injections and the torque being induced between both TMs along the $y$ axis.
The estimation of magnetic forces experienced by the test masses during the magnetic injections is performed by demodulating the $\Delta g$ measurements. Terms $\hat{F}_{DC,x}$, $\hat{F}_{1\omega,x}$ and $\hat{F}_{2\omega,x}$ previously defined in Section III.2 are estimated by applying a heterodyne demodulation at the corresponding frequencies and rescaling the amplitudes obtained by means of the mass of the TMs, $M_{\rm{TM}}=(1.928\pm 0.001)$ kg. Analogously, the same procedure can be extrapolated to the torque by using the moment of inertia of a cube with the side length of the TMs, $l_{\rm{TM}}=\rm(46.000\pm 0.005)$ mm.
@@
Since the magnetic field can not be directly measured in the test mass position we must refer to the magnetometers read-out for calibration. We measured the amplitude in the $+x$ magnetometer of each of the 20 injections of June $\rm 18^{th}$, 2016 for coil #$\#1$. To do so we demodulate the read-out at the injection frequency. By comparing these amplitudes to the ones predicted by Ampère’s law with origin at the coil location we found a systematic discrepancy of $11.85\pm 0.45\%$, with the predicted magnetic field being larger than the measurements from the magnetometers. This systematic discrepancy can have many origins. For the magnetic field prediction, we are assuming a perfect alignment of both the coil and magnetometer, as well as no tilts. In reality, during both manufacturing and integration of the magnetic diagnostics items (coils and magnetometers) there are unavoidable mechanical tolerances. Furthermore, the launch itself adds more uncertainty on the relative distance and tilts between the diagnostics items, being this a systematic error that is difficult to quantify a priori. All the effects combined can add up to the value reported. Since the magnetometer accuracy is 0.5%, we assume that the mismatch originates in the generation of the magnetic field by the coils and we apply this correction to all the calculated magnetic field values that intervene in all the equations derived in the previous section, redefining them for simplicity, i.e. $B_{x}^{AC}\equiv 0.8815\left\langle B_{x}^{AC}\right\rangle$.
@@ IV.1 Remanent magnetic moment
$\hat{F}_{1\omega,x}=\left[M_{r,x}+\frac{2\chi V}{\mu_{0}}B_{x}^{DC}\right]\nabla_{x}B_{x}^{AC}.$ (15)
@@
In order to evaluate this term, during the June $\rm 18^{th}$ run several injections with different AC field gradients were applied to the TMs. By doing so we can evaluate the term in brackets above at different values of the gradient $\nabla_{x}B_{x}^{AC}$. This is shown in Fig. 3 where we display how the $1\omega$ component of the force changes depending on the intensity of the injected AC magnetic field gradient, for fixed DC offsets. Each result in the plot represents an injection at 5 mHz with AC amplitudes applied to the coil of $I^{AC}=0.5,0.8,1.0,1.5\,{\rm mA}$.
As explained in Section III.2, by running the experiment at different DC levels we can further disentangle the dependencies of the parameters inside the brackets and obtain the estimate for the remanent magnetic moment. Table 1 gathers the linear fits to the results that we also show in Fig. 4. The offset parameter corresponds to the remanent magnetic moment of test mass #$\#1$ in the $x$ direction, $M_{r,x}=0.140\pm 0.138$ nAm2. The slope parameter is directly related to the magnetic susceptibility at $1\omega$. However, the values that were used for DC offsets of ±$\pm 0.1,0.2$ mA came from the injections of April $\rm 28^{th}$ which were performed at different frequencies (3 mHz), than the rest at 5 mHz. Thus, to obtain the value of the susceptibility at 5 mHz, we will use the fit from Fig. 4, but with only the DC values of ±$\pm 0.75,1.5$ mA giving a result of $\chi_{\rm{5mHz}}=(-3.3723\pm 0.0069)\times 10^{-5}$.
@@
Figure 4: Effective remanent magnetic moment plotted as a function of the injected DC magnetic field. The result of the fit, for an equation of the type y = mx + n, is m = $(-3.380\pm 0.027)\times 10^{-5}$ and n = $(0.140\pm 0.138)\times 10^{-9}$ Am2.
Analogously, to obtain $M_{r,y}$ and $M_{r,z}$, we demodulate the amplitudes of the torque measurements around the required angles and apply Eq. (13) for the respective injected magnetic fields. This way, we obtained $M_{r,y}=0.178\pm 0.025$ nAm2 and $M_{r,z}=0.095\pm 0.010$ nAm2 which we will need for the background estimations. Note the large uncertainty obtained for $M_{r,x}$ compared with $M_{r,y}$ and $M_{r,z}$. The difference is due to the different methods used to extract them. The former is calculated indirectly as the offset of the linear fit of the slopes of the force at $1\omega$ while the other two components of the remanent magnetic moment are calculated directly from the measured torques along $\eta$ and $\phi$. Rotations along $\theta$ would have allowed a more precise value for $M_{r,x}$ but the interferometric system is not sensitive along such axis as it is aligned with the laser beam. The results lead to a total remanent magnetic moment of: $|\textbf{M}_{r}|=(0.245\pm 0.081)$ nAm2.
@@ IV.2 Background magnetic field
The induction of forces in the test mass by means of the controlled injection of magnetic fields allow the determination not only of the test mass magnetic parameters but also of environment parameters that contribute to the magnetic force, which is the case of the background magnetic field in the test mass position. We emphasize here that this is the way to estimate of this parameter since the magnetometers are located too far away from the test masses to guarantee a precise estimate of this variable.
In order to do so we evaluate Eqs. (14) using the injections of June $\rm 18^{th}$, 2016. We express the information provided by these runs by displaying the DC component of the measured force as a function of the DC component of the applied magnetic fields. As predicted, we obtain the parabolas in Fig. 5 for different values of the AC amplitude from where we derive the parabola coefficients of Table 2.
@@
As derived from Eq. (14), the $A$ coefficient provides a direct estimate of the magnetic susceptibility
@@
which, in contrast with other alternative estimates, is not dependent of the injection modulation frequency. The value obtained using Eq. (16) is $\chi_{DC}=(-3.35\pm 0.15)\times 10^{-5}$.
With two remaining terms, $B$ and $C$, we can build a system of equations, being the two unknowns the parameters that define the background magnetic field in the test mass position, i.e. $B_{back.,x}$ and $\nabla_{x}B_{back.,x}$
@@
$B=\frac{M_{r,x}}{\kappa}+\frac{\chi V}{\mu_{0}}\left[\nabla_{x}B_{back.,x}+\frac{B_{back.,x}}{\kappa}\right].$ (18)
@@ IV.3 Magnetic Susceptibility
We have already estimated the test mass magnetic susceptibility as a by-product of the estimate of the remanent magnetic moment and the background magnetic field and field gradient.
@@
If we now plot the measured values of the magnetic susceptibility for TMs 1 and 2 along with the curve in Eq. (19), we obtain the plot shown in Fig. 6. We can see that all magnetic susceptibility results from Table 3 are compatible within their uncertainty ranges as predicted from Eq. (19), except for the one at 30 mHz. There were only three signal injected at 15 mHz, which had a low SNR and, also, showed an unexpected linear drift that had to be subtracted. We think this explains the systematics affecting the demodulation of its second harmonic at 30 mHz. With respect to the results for TM2, we computed them for completeness but we recall that coil 2 suffered a malfunction at the beginning of operations which could explain the systematic error observed in the derived parameters.
The results shown were obtained by heterodyne demodulation of the $\Delta g$ signal with a sinusoidal signal in-phase. However, if we made the latter be out of phase by $\pi/2$ one would expect the amplitudes measured by this method to be zero only if there were no imaginary component. Thus, the imaginary susceptibility can be obtained demodulating at $2\omega$, in quadrature. In the frequency regime that we are working in, this value is expected to be orders of magnitude smaller than the real susceptibility. The results that we obtain for the imaginary susceptibility at 10 mHz were consistent with zero, $\chi_{i}=(0.0\pm 1.8)\times 10^{-6}$. With the limits determined by the force sensitivity of the demodulation around the tenths of femtoNewtons. Thus, we can only confirm that the values of the imaginary susceptibility of the TMs are below $|\chi_{i}|<1.8\times 10^{-6}$ at 10 mHz. At the rest of frequencies the result gave a less precise upper bound for the imaginary susceptibility.
Finally, we can evaluate the prediction of our force model from Eq. (7) in comparison with the measured acceleration during injections by making use of the extracted magnetic parameters within this article. To do so, we have selected a single injection from all the ones of June $18^{\rm{th}}$, 2016 with a DC offset of 0.75 mA and an AC amplitude of 1.5 mA. We have used these values for the corresponding magnetic field DC and AC calculations together with the magnetic parameters found in the previous sections and substituted into Eqs. (8). The results can be seen in Fig. 7 where, together with the total model force, we can see plotted the different force contributions $F_{DC}$, $F_{1\omega}$ and $F_{2\omega}$. The predicted force and the data match with a residual difference between them of $(0.0\pm 1.9)\times 10^{-13}$ N. Due to the presence of a linear drift within the data originated from the desynchronization between the clocks of the LPF measurement systems, one for the diagnostics and a different one for the optical metrology system, we are able to observe a leftovers signal within the residual data.
@@ V Conclusions
The work in this paper presents the most detailed characterization of the magnetic-induced coupling on free-falling test masses in the context of gravitational wave detection in space. The results and method showed here were originated in the framework of the LISA Pathfinder mission. As a technology demonstrator, LPF represented a unique opportunity for an in-depth description of these effects, which can be directly transferred to LISA and future space-borne gravitational wave detectors.
In our results, we obtain the magnetic parameters defining the response of the free-falling test masses under an external magnetic field. According to our measurements, the test mass remanent magnetic moment does not show any privileged direction, which is in agreement with the hypothesis of an isotropic, diamagnetic test mass. We have found the value of its modulus to be: $\rm|\textbf{M}_{r}|=(0.245\pm 0.081)\,\rm{nAm}^{2}$, a result with much better precision than any other on-ground tests and also below the mission requirements of $\rm|\textbf{M}_{r}|\leq 10\,nAm^{2}$. The background magnetic field, $B_{back.,x}=414\pm 74$ nT, and its gradient, $\nabla_{x}B_{back.,x}=-7400\pm 2100$ nT/m, calculated at the location of the TMs are in agreement with the values predicted by  [16]. Moreover, the gradient is also below the worst case scenario predicted by  [19] which estimated a value of $\nabla_{x}B_{back.,x}=-11300$ nT/m. The TMs magnetic susceptibility results follow the model predicted by  [17] and they are also within the same magnitude when comparing the on-ground estimations for the DC magnetic susceptibility with values around $-2.5\times 10^{-5}$  [20] with our in-flight DC prediction for the magnetic susceptibility, $\chi_{DC}=(-3.35\pm 0.15)\times 10^{-5}$ and the most precisely characterized at 5 mHz of $\chi_{\rm 5mHz}=(-3.3723\pm 0.0669)\times 10^{-5}$. Finally, the value of the imaginary component of the magnetic susceptibility of the TMs has been set to an upper bound at 10 mHz such that: $|\chi_{i}|<1.8\times 10^{-6}$.
These in-flight results have been acquired with precisions not achieved before with on-ground measurements. They will be of great relevance for the design of future missions with similar technologies to the ones used in LPF.
@@ Appendix A Magnetic field calculation
When the magnetic field is created by an induction coil, such as the case in Fig. 1, the laws of Classical Magnetic Theory can be applied to obtain formulas which produce the values of the field and its gradient at any position in space. For slowly varying coil currents and short distances, radiative effects can be neglected.
@@ Appendix B Magnetic torque
$\begin{split}\text{N}_{\eta}=&\left\langle\sin(\omega t)\left(m_{r,z}B_{x}^{AC}-m_{r,x}B_{z}^{AC}\right)\right\rangle\\ &+\left\langle\left[r_{z}\Omega_{x}-r_{x}\Omega_{z}\right]\right\rangle,\end{split}$ (45)
@@
$\begin{split}\mathbf{\Omega}=&\left(\mathbf{m}_{r}\cdot\mathbf{\nabla}\right)\mathbf{B}^{AC}\sin(\omega t)\\ &+\dfrac{\chi}{\mu_{0}}\left(\dfrac{1-\cos(2\omega t)}{2}\right)\left(\mathbf{B}^{AC}\mathbf{\nabla}\right)\mathbf{B}^{AC},\end{split}$ (46)
@@
$\text{N}_{\eta}=\left\langle m_{r,z}B_{x}^{AC}\right\rangle\sin(\omega\,t).$ (47)
@@
$\text{N}_{\phi}=\left\langle-m_{r,y}B_{x}^{AC}\right\rangle\sin(\omega\,t).$ (48)
@@
$\begin{split}\mathbf{N}_{DC}=\left\langle\mathbf{m_{r}}\times\mathbf{B}_{0}\right\rangle,\end{split}$ (50a)
$\begin{split}\mathbf{N}_{1\omega}=\left\langle\mathbf{m_{r}}\times\mathbf{B}^{AC}\right\rangle\sin(\omega\,t).\end{split}$ (50b)